A stone is thrown downward at 30 m/s. What is its speed, [2]






a)after 1s b) after 2s

1
Expert's answer
2021-11-14T17:10:28-0500

v=u+at

a=9.81

u=30 m/s

a) t=1 s

"v=30 + 9.81 \times 1 = 39.81\; m/s"

b) t= 2 s

"v= 30 + 9.81 \times 2 = 49.62\; m/s"